# url]https://ntdxyg.github.io/

### v1

@@ Abstract
{highlights}
We define the pruning objectives for generative coding tasks, i.e., the KL divergence between the probability distribution of the pruned model and the original model, which are different from previous studies.
We propose Flab-Pruner pruning method with unified structure integrates vocabulary, layer, and FFN pruning, effectively reduces model parameters and performs well in generative coding tasks.
We introduce a customized code instruction tuning strategy for coding tasks to enhance the performance recovery efficiency of the pruned model.
We propose a new dataset CodeHarmony comprising 16K samples for generative coding tasks, with train/validation/test splits, which encompasses code generation, cot generation, and code output prediction tasks.
The pruned models exhibit significant improvements in storage, GPU usage, computational efficiency, and environmental impact, while maintaining good robustness.
@@ 1 Introduction
    Misaligned Pruning Objectives: Current pruning methods [44, 72, 29] primarily focus on layer-wise similarity metrics (e.g., angle distance, cosine similarity, and Taylor score between layers), which aim to preserve the model’s general language modeling capabilities while overlooking the specific requirements of downstream tasks.
@@
    Lack of Code-specific Post-tuning: Existing pruning methods rely on generic supervised fine-tuning on downstream datasets for performance recovery, without considering domain-specific adaptations for code-related tasks. This generic approach fails to leverage the unique characteristics and requirements of code generation tasks, potentially limiting the effectiveness of the post-pruning recovery process.
@@ 3 Our Method
The workflow of Flab-Pruner is shown in Figure 1. There are three major pruning steps, i.e., vocab pruning, layer pruning and FFN pruning. For a given sense model, we first perform vocab pruning, then layer pruning, and finally FFN pruning.
@@ 13                   $\mathcal{D}_{\text{correct}}\leftarrow\mathcal{D}_{\text{correct}}\cup\{sample\}$;
17 Function FindBestLayerToPrune($\mathcal{M}_{\text{pru}},H^{(L)}_{\text{original}}$):
18       $best\_score\leftarrow-\infty$;
@@ 3.5 Performance Recovery
After pruning some model parameters, it is crucial to implement effective strategies to recover the model’s performance. The following pseudocode describes our proposed Performance Recovery Algorithm.
@@ 5.1 RQ1: Performance Comparison
One of the primary objectives of this study is to evaluate the effectiveness of Flab-Pruner in reducing model size while preserving the performance. To this end, we design a series of experiments to compare the performance of Flab-Pruner against other structured pruning methods as well as the original dense models. We consider five representative pruning methods, including ShortGPT, UIDL, Linearity, SLEB and LLM-pruner as the baselines. Each of these methods has a unique focus and explores different aspects of model pruning. For each baseline, we control it to have a similar number of model parameters as Flab-Pruner after pruning to ensure fairness.
@@
(2) Compared with other pruning objectives: As shown in Figure LABEL:fig:layer, we compare the effect of pruning different number of layers on the model performance and find that the more layers are pruned, the more the model performance decreases. Furthermore, we conduct an in-depth analysis of our proposed pruning objectives in Layer Pruning and find that our pruning objectives can maintain higher performance compared to other objectives proposed by baselines.
@@
(4) Ablation Study: To further understand the individual contributions of different components within our approach, we conduct an ablation study to evaluate the impact of each component on the overall performance of the pruned models. As shown in Table 3, the results of the ablation study indicate that the vocab pruning strategy can not have a significant impact on the performance of the pruned models. The FFN pruning strategy, on the other hand, shows a slight decrease in performance compared to the dense model.
@@ 5.2 RQ2: Efficiency Analysis
Summary of RQ2 The efficiency comparison underscores the practical advantages of Flab-Pruner in resource-constrained settings, providing a sustainable solution for deploying powerful code intelligence models with reduced computational overhead.
@@ 5.3 RQ3: Robustness Analysis
6 Threats to Validity
@@ 8 Conclusion and Future Work
In future work, we plan to explore more Code LLMs and evaluate the effectiveness of Flab-Pruner on a broader range of generative coding tasks. We will release more pruned models on the open source platform to facilitate the deployment of pruned models in real-world software engineering scenarios. Moreover, the current work focuses on Python code. We plan to expand Flab-Pruner to support a broader range of programming languages in the future.

### v2

@@ 1 Introduction
    Misaligned Pruning Objective: Current pruning methods [45, 73, 29] primarily focus on layer-wise similarity metrics (e.g., angle distance, cosine similarity, and Taylor score between layers), which aim to preserve the model’s general language modeling capabilities while overlooking the specific requirements of downstream tasks.
@@
    Insufficient Code-specific Post-tuning Method: Existing pruning methods rely on generic supervised fine-tuning on downstream datasets for performance recovery, without considering domain-specific adaptations for code-related tasks. This generic approach fails to leverage the unique characteristics and requirements of code generation tasks, potentially limiting the effectiveness of the post-pruning recovery process.
@@ 3 Our Method
The workflow of Flab-Pruner is shown in Figure 1. There are three major pruning steps, i.e., vocab pruning, layer pruning and FFN pruning. For a given dense model, we first perform vocab pruning, then layer pruning, and finally FFN pruning.
@@ 13                   $\mathcal{D}_{\text{correct}}\leftarrow\mathcal{D}_{\text{correct}}\cup\{sample\}$;
17 Function FindBestLayerToPrune($\mathcal{M}_{\text{pru}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}P_{\text{original}}}$):
18       $best\_score\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\infty}$;
@@ 3.5 Performance Recovery
After pruning some model parameters, it is crucial to implement effective strategies to recover the model’s performance. For performance recovery, we fine-tune the pruned model on the training split of our CodeHarmony dataset (detailed in Section 4.3), which contains 15,800 samples for code generation and CoT tasks and 100,706 samples for output prediction tasks. The following pseudocode describes our proposed Performance Recovery Algorithm.
@@ 5.1 RQ1: Performance Comparison
One of the primary objectives of this study is to evaluate the effectiveness of Flab-Pruner in reducing model size while preserving the performance. To this end, we design a series of experiments to compare the performance of Flab-Pruner against other structured pruning methods as well as the original dense models. We consider five representative pruning methods, including ShortGPT, UIDL, Linearity, SLEB and LLM-pruner as the baselines. Each of these methods has a unique focus and explores different aspects of model pruning. For each baseline, we control the pruning ratio to be 20% (slightly lower than our method’s 22%) to ensure a fair comparison across all methods.
@@
(2) Compared with other pruning objectives: As shown in Figure LABEL:fig:layer, we compare the effect of pruning different number of layers on the model performance and find that the more layers are pruned, the more the model performance decreases. Furthermore, we conduct an in-depth analysis of our proposed pruning objectives in Layer Pruning. We conducted comprehensive comparisons between our KL divergence-based approach and several alternative pruning objectives. When examining the impact of the progressively increasing number of pruned layers, we observe that with minimal pruning (1-2 layers), all methods maintain performance comparable to the original model. However, as more layers are removed, significant differences emerge between pruning objectives. The angular distance-based method (UIDL) deteriorates most dramatically, with pass@1 metrics dropping from above 60% to 0% when pruning beyond a certain threshold. Both Taylor formula-based gradients (LLM-Pruner) and cosine similarity-based methods (ShortGPT) show substantial performance degradation as pruning increases, while our KL divergence approach maintains higher performance with a more gradual decline. When pruning ratios become extremely high, all methods eventually experience severe performance drops, though our approach consistently degrades more gracefully. These results demonstrate that KL divergence captures layer importance for code intelligence tasks more effectively than both gradient-based and representation similarity-based methods.
@@
(4) Ablation Study: To further understand the individual contributions of different components within our approach, we conduct an ablation study to evaluate the impact of each component on the overall performance of the pruned models. Table 3 demonstrates that vocabulary pruning contributes minimally to performance changes in the pruned models, as evidenced by our ablation analysis. The FFN pruning strategy, on the other hand, shows a slight decrease in performance compared to the dense model.
Furthermore, we conducted a comparative analysis between our heuristic rule-based methods and a neuron importance evaluation approach using Fisher Information Matrix (FIM [43]), which is shown in Table 3. The FIM-based approach achieves similar performance compared to our Top-K method but with a much higher computational cost. Our proposed method requires approximately one minutes while the FIM-based approach requires approximately 30 minutes. In addition, the FIM-based approach requires additional forward and backward passes through the network for each batch of data.
Based on these results, we validate our design choice and demonstrate that, while contribution-based method may offer theoretical advantages, our heuristic rule-based approach provides comparable performance with significantly reduced computational requirements.
@@ 5.2 RQ2: Efficiency Analysis
End-to-End Efficiency Analysis: To provide a comprehensive view of the efficiency gains across the entire model lifecycle, we analyze the one-off computational cost of pruning and fine-tuning versus the recurring benefits during inference. This analysis addresses the complete efficiency landscape from model compression to deployment.
Table 6: Comprehensive efficiency analysis across different stages of Flab-Pruner deployment in One-time Cost and Inference Efficiency.
Stage Time (hours) GPU Memory (GB) FLOPs (T)
Dense - 13.55 7.04
+ Vocabulary Pruning 0.02 12.15 6.80
+ Layer Pruning 0.3 11.10 5.98
+ FFN Pruning 0.04 10.72 5.64
+ Performance Recovery 5.5 10.72 5.64
Flab-Pruner 5.86 10.72 5.64
As shown in Table 6, the entire pruning and recovery process takes approximately 6 hours on RTX 3090. While this represents a high initial investment, our analysis shows that these costs are quickly amortized through inference efficiency gains.
For the computational efficiency analysis, we need to consider the one-off cost of pruning versus the recurring savings during inference. In terms of FLOPs, each inference with the dense model requires 7.04T FLOPs, while the pruned model requires only 5.64T FLOPs, saving 1.4T FLOPs (19.9%) per inference. Considering the dense model’s approximate throughput of 7.04T FLOPs during our experiments, the 6-hour pruning and fine-tuning process consumes approximately $6\times 3600\times 7.04=152,064$T FLOPs (we assume that the frequency of calculation is 1 time/second). To determine the break-even point, we calculate:
[…]
Namely, the computational investment is redeemed after approximately 108,617 inference runs.
For real-world deployments where models are typically invoked millions of times especially in production environments or as part of developer tools and code assistants, this initial cost is readily justified. For example, a deployment serving 10,000 inference requests daily would redeem the initial pruning investment in approximately 11 days, after which all efficiency gains represent net computational savings.
Summary of RQ2 The efficiency comparison underscores the practical advantages of Flab-Pruner in resource-constrained settings, providing a sustainable solution for deploying powerful code intelligence models with reduced computational overhead. When considering the full model lifecycle, the initial pruning costs are quickly amortized, resulting in significant efficiency gains for real-world deployments.
@@ 5.3 RQ3: Robustness Analysis
6 Discussion
6.1 Evaluation on BigCodeBench
To ensure a thorough evaluation, we conduct additional experiments on BigCodeBench [82], a more diverse and challenging benchmark designed to evaluate code models across various dimensions of programming tasks.
BigCodeBench consists of 1,140 function-level tasks specifically designed to challenge language models in following instructions and composing multiple function calls from 139 different libraries as tools. Unlike HumanEval and OpenEval, BigCodeBench provides a more rigorous evaluation framework with comprehensive test coverage, complex user-oriented instructions, open-ended problem solving and tool composition.
Table 8: Performance comparison on BigCodeBench (Pass@1 %)
Model Dense Flab-Pruner
w/o PT w/ PT
CodeQwen 39.60 36.20 39.80
NxCode 39.60 36.10 40.00
CodeSlerp 39.80 37.10 39.50
As shown in Table 8, our findings on BigCodeBench align with our main results:
  • Without PT, the pruned models maintain approximately 91-93% of the dense models’ performance, demonstrating reasonable retention of capabilities despite parameter reduction.
  • With PT, the pruned models achieve performance comparable to or slightly better than the dense models, with CodeQwen and NxCode showing small improvements of 0.2-0.4% in Pass@1 scores, while CodeSlerp experiences a negligible decrease of 0.3%.
  • The models’ ability to compose function calls from different libraries remains intact after pruning, indicating that Flab-Pruner preserves the models’ semantic understanding and tool utilization capabilities.
The consistent performance patterns across this more challenging benchmark corroborate our findings from HumanEval and OpenEval, demonstrating that Flab-Pruner maintains its effectiveness even on tasks requiring complex reasoning and tool composition. It also demonstrates the generalizability of our approach across different task complexities and problem-solving scenarios.
6.2 Impact between Strategies
To address the question of how our three pruning strategies (vocabulary pruning, layer pruning and FFN pruning) interact with each other, we conduct a detailed ablation analysis of their individual and combined effects, which is performed on the CodeQwen model. Table 9 presents the results on the HumanEval benchmark and the parameter reduction achieved in each configuration.
Table 9: Impact of Pass@1 on different pruning strategy combinations on the CodeQwen model (w/o PT)
Pruning Strategy HumanEval OpenEval CodeHarmony
None (Dense) 77.44 42.13 60.78
Vocab only 77.44 42.70 62.09
Layer only 75.00 37.64 64.05
FFN only 76.54 39.80 66.00
Vocab + Layer 77.20 42.13 64.92
Vocab + FFN 74.39 37.64 64.05
Layer + FFN 76.24 36.52 66.83
All three (Flab-Pruner) 76.22 36.52 66.63
Our findings reveal several key insights:
  • Individual Impact: Vocabulary pruning demonstrates the least performance degradation, maintaining full performance on HumanEval (77.44%) and even slightly improving on OpenEval (42.70% vs. 42.13%). In contrast, layer pruning shows the most significant impact on HumanEval (dropping to 75.00%) and OpenEval (dropping to 37.64%). Interestingly, all three individual strategies actually improve performance on CodeHarmony, with FFN pruning providing the largest gain (66.00% vs. 60.78%).
  • Pairwise Combinations: The combination of vocabulary and layer pruning shows a synergistic effect, with minimal impact on HumanEval (77.20%) and no degradation on OpenEval (42.13%). However, when FFN pruning is combined with either vocabulary or layer pruning, we observe more significant performance drops on HumanEval and OpenEval. For CodeHarmony, all pairwise combinations yield substantial improvements, with layer+FFN pruning showing the highest gain (66.83%).
  • Complementary Effects: When all three strategies are combined (Flab-Pruner), we observe a balanced trade-off. The performance on HumanEval (76.22%) and OpenEval (36.52%) shows moderate decreases, while CodeHarmony performance improves substantially (66.63%). This suggests that the strategies have task-dependent interactions, with some benchmark tasks being more sensitive to specific types of pruning than others.
These findings indicate that the three pruning strategies target different aspects of the model’s parameter space and have complementary effects. While each strategy contributes to parameter reduction, their careful combination, followed by effective performance recovery, is essential for maintaining or even enhancing model capabilities across diverse coding tasks.
6.3 Threats to Validity
@@ 8 Conclusion and Future Work
In future work, we shall develop more efficient model pruning methods to reduce the computational overhead of our current approaches. Potential directions include cherry data selection techniques [32] to select a smaller dataset for evaluation and transfer-based methods that leverage importance patterns from previously pruned models to new architectures.
Additionally, we plan to explore more Code LLMs and evaluate the effectiveness of Flab-Pruner on a broader range of generative coding tasks. We will release more pruned models on the open source platform to facilitate the deployment of pruned models in real-world software engineering scenarios. Moreover, the current work focuses on Python code. We plan to expand Flab-Pruner to support a broader range of programming languages in the future.